OpenStudy (anonymous):

what is radical 250?

OpenStudy (anonymous):

simplify please

OpenStudy (anonymous):

5rad10

OpenStudy (anonymous):

rad 25 x 10 5rad10

OpenStudy (anonymous):

\[\sqrt{250}=\sqrt{25\times 10}=\sqrt{25}\times \sqrt{10}=5\sqrt{10}\]